OpenStudy (anonymous):

Express the complex number in trigonometric form. -2i 2(cos 180° + i sin 180°) 2(cos 90° + i sin 90°) 2(cos 270° + i sin 270°) 2(cos 0° + i sin 0°)

OpenStudy (gorv):

a+bi= 0+2i

OpenStudy (gorv):

how do we calculate the angle ??

OpenStudy (gorv):

also the value of r

OpenStudy (gorv):

\[r=\sqrt{a^2+b^2}\]

OpenStudy (gorv):

now you tell me abt how we cal angle??

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

use the formula you just said, right? If r is the angle then it is 4

OpenStudy (gorv):

r is not the angle

OpenStudy (gorv):

a+ib in trigonometric form =\[r*(\cos \theta + i \sin \theta)\]

OpenStudy (anonymous):

i don't know how to find the angle

OpenStudy (gorv):

\[\theta=\tan^{-1}\frac{ imaginary -part }{ real-part}\]

OpenStudy (gorv):

can u do it now ??

OpenStudy (gorv):

imaginary part=-2 real part= 0

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

63.434

OpenStudy (anonymous):

Wait, you can't divide by 0

OpenStudy (gorv):

anything divide by zero =???

OpenStudy (anonymous):

undifined? 0?

OpenStudy (anonymous):

undefined*

OpenStudy (gorv):

when tan becomes undefined???

OpenStudy (anonymous):

cos(theta) = 0?

OpenStudy (gorv):

when cos gives 0 ???

OpenStudy (gorv):

i mean for what value of theta cos becomes 0??

OpenStudy (anonymous):

90 and 270

OpenStudy (gorv):

well our imaginary part is negative so when there is negative in tan it will give negative output

OpenStudy (anonymous):

so 270

OpenStudy (gorv):

yes :)

OpenStudy (gorv):

now trigo form = r(cos theta + i sin theta )

OpenStudy (gorv):

tell me what you got ???

OpenStudy (anonymous):

2(cos 270° + i sin 270°)

OpenStudy (gorv):

how to calculate r you already know

OpenStudy (gorv):

yes

OpenStudy (gorv):

cos 270=??

OpenStudy (anonymous):

0

OpenStudy (gorv):

sin 270=???

OpenStudy (gorv):

you already got the ans

OpenStudy (anonymous):

-1

OpenStudy (gorv):

this is for back conversion to a+i b form

OpenStudy (gorv):

2(0+i(-1))=-2i so we are done :)

OpenStudy (gorv):

hope you wont face any difficulty in trigo or a+ib form conversion